# by Laurent Busé, Thomas Dedieu

### v2

@@ Abstract
For a given $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, the polynomial $\Phi_{p^{\prime},p^{\prime\prime}}$ splits as the product of $\mu$ linear forms $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$, $i=1,\ldots,\mu$, that all vanish along the line $\langle p^{\prime},p^{\prime\prime}\rangle$. Although this splitting may not exist globally over $\mathcal{T}_{S}$ (in other words the factors $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ may not be polynomials in $(p^{\prime},p^{\prime\prime})$), there exists a global splitting over a finite cover $\tilde{\mathcal{T}_{S}}$ of $\mathcal{T}_{S}$444 let us briefly recap how such a finite cover may be constructed. The family of planes defined over $\mathcal{T}_{S}$ by $\Phi_{p^{\prime},p^{\prime\prime}}$ corresponds to a subscheme $\mathcal{P}$ of $\mathcal{T}_{S}\times\check{\mathbf{P}^{3}}$ with the first projection inducing a degree $\mu$ map $\mathcal{P}\to\mathcal{T}_{S}$. Our goal is to show that there exists a base change $\tilde{\mathcal{T}_{S}}\to\mathcal{T}_{S}$ such that $\mathcal{P}\times_{\mathcal{T}_{S}}\tilde{\mathcal{T}_{S}}$ splits as $\mu$ copies of $\tilde{\mathcal{T}_{S}}$. As a first step, we consider the following (somewhat absurd non-sensical) base change. ${\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{S}}$ The diagonal $\Delta$ provides a section of the degree $\mu$ map $\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}\to\mathcal{P}$, hence $\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}$ splits as $\Delta+\mathcal{P}^{\prime}$ where $\mathcal{P}^{\prime}$ comes with a degree $\mu-1$ map to $\mathcal{P}$. Then, by induction on $\mu$, one may construct a finite cover $\tilde{\mathcal{P}}\to\mathcal{P}$ such that $\mathcal{P}^{\prime}\times_{\mathcal{P}}\tilde{\mathcal{P}}$ splits as $\mu-1$ copies of $\tilde{\mathcal{P}}$, hence $\mathcal{P}\times_{\mathcal{T}_{S}}\tilde{\mathcal{P}}$ splits as $\mu$ copies of $\tilde{\mathcal{P}}$. Thus $\tilde{\mathcal{P}}$ is suitable as the $\tilde{\mathcal{T}_{S}}$ we were looking for. We point out the (non-accidental) similarity of this construction with that of a splitting field for a given polynomial $P$ in one indeterminate $X$. The first base change in our construction is the analogue of the field extension $\mathbf{k}[X]/(P)$. Also note that, in general, the map $\tilde{\mathcal{T}_{S}}\to\mathcal{T}_{S}$ will have degree $\mu!$. (in other words the factors $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ have their coefficients in a finite extension of $\mathbf{k}[p^{\prime},p^{\prime\prime}]$). Then, by the multiplicativity property, see [15, § 5.7, p. 154], the resultant $R$ splits as the product of the $\mu$ resultants $R_{i}=\mathrm{Res}(T_{p^{\prime}},\Phi^{i}_{p^{\prime},p^{\prime\prime}},L^{\prime},L^{\prime\prime})$. By Lemma (3.3), for all $i$ the resultant $R_{i}$ vanishes wherever $D_{L^{\prime},L^{\prime\prime}}$ vanishes on $\mathcal{T}_{S}$, or rather on $\tilde{\mathcal{T}_{S}}$, as both forms $T_{p^{\prime}}$ and $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ vanish along the line $\langle p^{\prime},p^{\prime\prime}\rangle$. By Lemma (3.4), this implies that $D_{L^{\prime},L^{\prime\prime}}$ divides $R_{i}$ as a function of $(p^{\prime},p^{\prime\prime})\in\tilde{\mathcal{T}_{S}}$. Eventually, the upshot is that $(D_{L^{\prime},L^{\prime\prime}})^{\mu}$ divides $R$, and since both $D_{L^{\prime},L^{\prime\prime}}$ and $R$ are well-defined as functions of $(p^{\prime},p^{\prime\prime})\in{\mathcal{T}_{S}}$, so is the quotient. Therefore, there exists a multi-homogeneous polynomial $R^{\prime}$ in the sets of variables $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$ such that after restriction to $\check{\mathbf{P}}^{3}\times\check{\mathbf{P}}^{3}\times\mathcal{T}_{S}$ one has $R=(D_{L^{\prime},L^{\prime\prime}})^{\mu}R^{\prime}$. Computing degrees, one finds that $R^{\prime}$ has multi-degree $(0,0,\mu(d-2)+\lambda,0)$ in $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$, which proves the theorem. ∎
@@
We recall that $\mathbf{k}$ is assumed to be an algebraically closed field of characteristic 00. We let $S\subseteq\mathbf{P}^{3}$ be a smooth surface of degree $d$, defined by a homogeneous polynomial $F(X,Y,Z,W)$. We consider generic $4$-tuples $\hat{p}^{\prime}=(X^{\prime},Y^{\prime},Z^{\prime},W^{\prime})$, $\hat{p}^{\prime\prime}=(X^{\prime\prime},Y^{\prime\prime},Z^{\prime\prime},W^{\prime\prime})$, $\hat{p}=(X,Y,Z,W)$, and call $p^{\prime},p^{\prime\prime},p$ the corresponding points in $\mathbf{P}^{3}$ (beware the unusual distribution of the prime decorations). The choice of $\hat{p}^{\prime},\hat{p}^{\prime\prime},\hat{p}$ defines a system of homogeneous coordinates $(\alpha:\beta:\gamma)$ on the plane $\langle p^{\prime},p^{\prime\prime},p\rangle$ generated by $p^{\prime}$, $p^{\prime\prime}$ and $p$. We consider
@@
The coefficient of $f(\alpha,\beta,\gamma)$ in $\alpha^{a}\beta^{b}\gamma^{c}$ is a polynomial of degree $a$ in the set of indeterminates $\hat{p}^{\prime}$, as can be seen on the Taylor–Newton expansion, the first terms of which are given in Display (3.1.1). Therefore, by Proposition (2.6.1), the discriminant $\mathrm{Disc}(f)$ is homogeneous of degree $d(d-1)^{2}$ in $\hat{p}^{\prime}$, and the dual surface $S^{\vee}$ has degree $d^{\vee}=d(d-1)^{2}$ in $\smash{\check{\mathbf{P}}^{3}}$.
@@
[…]
@@
Salmon’s idea is that if the tangent plane $\mathbf{T}_{p^{\prime}}S$ is also tangent to $S$ at some point different from $p^{\prime}$, then it should appear with multiplicity $>2$ in $(p^{\prime},p^{\prime\prime})^{\perp}\cap S^{\vee}$; equivalently, $\mathrm{Disc}(f)$ should be divisible by $T^{3}$ in the above setup.
@@
The polynomial $A$ has degree 00 in $p$, therefore it is divisible by $T$ if and only if it is identically zero. By definition $A=\mathrm{D}^{\hat{p}^{\prime\prime 2}}f(\hat{p}^{\prime})$, so its vanishing is equivalent to the point $p^{\prime\prime}$ being on the polar quadric $\mathrm{D}_{{p^{\prime}}^{2}}S$ of $S$ at $p^{\prime}$. Since $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, this in turn is equivalent to the line $\langle p^{\prime},p^{\prime\prime}\rangle$ being one of the two inflectional tangents of $S$ at $p^{\prime}$ (i.e., the two lines $\Lambda$ having intersection multiplicity with $S$ at $p^{\prime}$ at least $3$; in symbols, $i(\langle p^{\prime},p^{\prime\prime}\rangle,S)_{p^{\prime}}\geqslant 3$).
@@
The polynomial $B^{2}-AC$ is tri-homogeneous of tri-degree $(2(d-2),2,2)$ in the variables $(p^{\prime},p^{\prime\prime},p)$. By Theorem (3.2), which have seen in Paragraph (3.7) above can be applied in our situation, there exists a homogeneous polynomial $H$ of degree $4(d-2)$ in the variable $p^{\prime}$, with constant coefficients, such that for fixed $p^{\prime}\in S$ and $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, $T$ divides $B^{2}-AC$ as polynomials in $p$ if and only if $H(p^{\prime})=0$.
@@
The discriminant in Display (3.4.2) is tri-homogeneous in the variables $(p^{\prime},p^{\prime\prime},p)$ with respect to the standard grading, of degree $d^{\vee}=d(d-1)^{2}$ with respect to all three variables, as we have observed above. This implies that also the homogeneous piece of (3.4.2) of lowest degree with respect to the Zariski grading is tri-homogeneous of tri-degree $(d^{\vee},d^{\vee},d^{\vee})$ for the standard grading. Moreover, $T$, $A$, and $B^{2}-AC$ have respective tri-degrees $(d-1,0,1)$, $(d-2,2,0)$, and $(2(d-2),2,2)$ in $(p^{\prime},p^{\prime\prime},p)$. Thus, computing degrees, one finds that $\Phi$ has tri-degree $(\lambda,\mu,\mu)$, with
@@
Since $\Phi$ is $\operatorname{redDisc}_{d}^{2}(f)$ with respect to the indeterminate $\alpha$, and the coefficient of $f(\alpha,\beta,\gamma)$ in $\alpha^{a}\beta^{b}\gamma^{c}$ is a polynomial of degree $a$ in the indeterminates $p^{\prime}$, the fact that $\Phi$ is homogeneous in $p^{\prime}$ and its degree follow directly from Proposition (2.13.1); to wit, in our situation the degree in Proposition (2.13.1) reads
@@
The homogeneity and degrees in $p$ and $p^{\prime\prime}$ however do not directly follow from Proposition (2.13), as for instance the coefficients in $A\beta^{2}+2B\beta\gamma+C\gamma^{2}$ do not all have the same weights in $p$ or $p^{\prime\prime}$. It is noticeable that $T^{2}A(B^{2}-AC)^{2}$ is homogeneous in $p$ and $p^{\prime\prime}$ because the weights of $T^{2}$ and $A$ balance each other; this is a very elementary reason why $T$ “should” appear with a square in Formula (3.4.2). The degree of $\operatorname{redDisc}_{d}^{2}(f)$ with the weights as in Proposition (2.13.2) is
@@
It has to be recognized that the proof given above of the latter statement rests on the not fully justified fact that $T$ divides $\Phi=\operatorname{redDisc}(f)$ if and only if the tangent plane $\mathbf{T}_{p^{\prime}}S$ is also tangent to $S$ at some additional point. The “if” part is established, see, e.g., [A, Theorem **] due to Dimca, but what is really needed is the “only if” part. Evidence in favour of that assertion is provided by the irreducibility of the reduced discriminant $\Phi$ (which, as already mentioned, follows from the fact that the reduced discriminant is a toric discriminant; see also [4]).
@@
[…]
Let $V=\mathrm{V}(F)$ be a smooth hypersurface of degree $d$ in $\mathbf{P}^{n}$, and consider $n$ points $p_{1},\ldots,p_{n-1},p$ (or rather tuples $\hat{p}_{1},\ldots,\hat{p}_{n-1},\hat{p}\in\mathbf{k}^{n+1}$). The points $p_{1},\ldots,p_{n-1}$ (if in linear general position) define a pencil of hyperplanes in $\mathbf{P}^{n}$, and for all $p$ (not in $\langle p_{1},\ldots,p_{n-1}\rangle$), the tuples $\hat{p}_{1},\ldots,\hat{p}_{n-1},\hat{p}$ define a system of homogeneous coordinates $(\alpha_{1}:\ldots:\alpha_{n-1}:\alpha)$ on the member $H_{p}$ of the pencil determined by $p$, i.e., $\langle p_{1},\ldots,p_{n-1},p\rangle$. In this system of coordinates, the hyperplane section $H_{p}\cap V$ is the zero locus of the polynomial
@@
[…]
@@
with $\bar{f}_{2}(\alpha_{2},\ldots,\alpha_{n})=f_{2}(\alpha_{2},\ldots,\alpha_{n},0)$ and $\Phi=\operatorname{redDisc}_{d}^{2}(f)$. Our task is to analyze the divisibilities of $\mathrm{Disc}(f_{2})$ and $\Phi$ by $T$ ($\bar{f}_{2}$ has degree 00 in $p$, hence $T$ divides $\mathrm{Disc}(\bar{f}_{2})$ if and only if $\mathrm{Disc}(\bar{f}_{2})=0$).
Theorem (3.2) (Salmon’s elimination trick) generalizes as follows. Given $n-1$ linear forms $L_{1},\ldots,L_{n-1}\in\mathbf{P}^{n}$, the line $\mathrm{V}(L_{1},\ldots,L_{n-1})$ intersects the $(n-2)$-dimensional linear space $\langle p_{1},\ldots,p_{n-1}\rangle$ if and only if
[…]
@@
and consider a multihomogeneous polynomial $\Phi_{p_{1},\ldots,p_{n-1}}(p)$ of multidegree $(\lambda,\mu,\ldots,\mu)$ in the sets of variables $p_{1},\ldots,p_{n-1},p$, such that for all $(p_{1},\ldots,p_{n-1})\in\mathcal{T}_{V}$, the hypersurface $\mathrm{V}(\Phi_{p_{1},\ldots,p_{n-1}})$ consists of $\mu$ hyperplanes, counted with multiplicities, all containing the $(n-2)$-plane $\langle p_{1},\ldots,p_{n-1}\rangle$.
@@
in $p_{1}$, and $p_{2},\ldots,p_{n-1},p$, respectively (note that the former degree is divisible by $d-2$). As in Paragraph (3.9), the degree in $p_{1}$ of $\operatorname{redDisc}_{d}^{2}(f)$ is given by Proposition (2.13.1), while the degree with respect to the grading of Proposition (2.13.2) is equally distributed between the degrees in $p_{2},\ldots,p_{n-1},p$ respectively.
@@
Let $V$ be a smooth degree $d$ hypersurface in $\mathbf{P}^{n}$, $n>1$. The number of hyperplanes bitangent to $V$ passing through $n-2$ fixed general points in $\mathbf{P}^{n}$ is
@@
In this section, we elaborate on the computation by Salmon of the number of bitangent lines to a surface in $\mathbf{P}^{3}$ passing through a general point. We give his proof in Section 4.1, and a generalization to hypersurfaces in $\mathbf{P}^{n}$ in the next Section 4.2. These proofs only involve the reduced discriminant of polynomials in two indeterminates, which is easily expressed as a plain discriminant, see Example (2.9). We give however another generalization in Section 4.3, for which it is necessary to consider reduced discriminants of polynomials with an arbitrary number of indeterminates. The two statements of Sections 4.1 and 4.2 are instances of a vast generalization by Fehér and Juhász [10], which is described in [B, Theorem **].
@@
There exists a polynomial $R$, bihomogeneous in $p^{\prime},p^{\prime\prime}$ with respective degrees $(d-2)(d-3)$ and $(d+2)(d-3)$, such that the locus of points $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$ such that the line $\langle p^{\prime},p^{\prime\prime}\rangle$ is bitangent to $S$ is cut out on $\mathcal{T}_{S}$ by $V(R)$.
@@
(this is an abuse of notation: actually one should consider two liftings $\hat{p}^{\prime},\hat{p}^{\prime\prime}\in\mathbf{k}^{4}$ of $p^{\prime}$ and $p^{\prime\prime}$ respectively, and $F(\alpha\hat{p}^{\prime}+\beta\hat{p}^{\prime\prime})$); the last equality comes from the fact that $\langle p^{\prime},p^{\prime\prime}\rangle$ is tangent to $S$ at $p^{\prime}$. Let $\vphantom{f}{}^{\flat}\kern-1.00006ptf(\alpha,\beta)$ be the homogeneous polynomial of degree $d-2$ between parentheses at the right-hand-side of (4.2.1). The line $\langle p^{\prime},p^{\prime\prime}\rangle$ is a bitangent to $S$ if and only if $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ has a multiple root, so the polynomial $R$ we are looking for is merely the discriminant of $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ which is none other, up to sign, than the reduced discriminant of $f$, as noticed in (2.9).

### v3

@@
For a given $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, the polynomial $\Phi_{p^{\prime},p^{\prime\prime}}$ splits as the product of $\mu$ linear forms $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$, $i=1,\ldots,\mu$, that all vanish along the line $\langle p^{\prime},p^{\prime\prime}\rangle$. Although this splitting may not exist globally over $\mathcal{T}_{S}$ (in other words the factors $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ may not be polynomials in $(p^{\prime},p^{\prime\prime})$), there exists a global splitting over a finite cover $\tilde{\mathcal{T}_{S}}$ of $\mathcal{T}_{S}$666let us briefly recap how such a finite cover may be constructed. The family of planes defined over $\mathcal{T}_{S}$ by $\Phi_{p^{\prime},p^{\prime\prime}}$ corresponds to a subscheme $\mathcal{P}$ of $\mathcal{T}_{S}\times\check{\mathbf{P}^{3}}$ with the first projection inducing a degree $\mu$ map $\mathcal{P}\to\mathcal{T}_{S}$. Our goal is to show that there exists a base change $\tilde{\mathcal{T}_{S}}\to\mathcal{T}_{S}$ such that $\mathcal{P}\times_{\mathcal{T}_{S}}\tilde{\mathcal{T}_{S}}$ splits as $\mu$ copies of $\tilde{\mathcal{T}_{S}}$. As a first step, we consider the following (somewhat absurd non-sensical) base change. ${\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{S}}$ The diagonal $\Delta$ provides a section of the degree $\mu$ map $\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}\to\mathcal{P}$, hence $\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}$ splits as $\Delta+\mathcal{P}^{\prime}$ where $\mathcal{P}^{\prime}$ comes with a degree $\mu-1$ map to $\mathcal{P}$. Then, by induction on $\mu$, one may construct a finite cover $\tilde{\mathcal{P}}\to\mathcal{P}$ such that $\mathcal{P}^{\prime}\times_{\mathcal{P}}\tilde{\mathcal{P}}$ splits as $\mu-1$ copies of $\tilde{\mathcal{P}}$, hence $\mathcal{P}\times_{\mathcal{T}_{S}}\tilde{\mathcal{P}}$ splits as $\mu$ copies of $\tilde{\mathcal{P}}$. Thus $\tilde{\mathcal{P}}$ is suitable as the $\tilde{\mathcal{T}_{S}}$ we were looking for. We point out the (non-accidental) similarity of this construction with that of a splitting field for a given polynomial $P$ in one indeterminate $X$. The first base change in our construction is the analogue of the field extension $\mathbf{k}[X]/(P)$. Also note that, in general, the map $\tilde{\mathcal{T}_{S}}\to\mathcal{T}_{S}$ will have degree $\mu!$. (in other words the factors $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ have their coefficients in a finite extension of $\mathbf{k}[p^{\prime},p^{\prime\prime}]$). Then, by the multiplicativity property, see [15, § 5.7, p. 154], the resultant $R$ splits as the product of the $\mu$ resultants $R_{i}=\mathrm{Res}(T_{p^{\prime}},\Phi^{i}_{p^{\prime},p^{\prime\prime}},L^{\prime},L^{\prime\prime})$. By Lemma (3.3), for all $i$ the resultant $R_{i}$ vanishes wherever $D_{L^{\prime},L^{\prime\prime}}$ vanishes on $\mathcal{T}_{S}$, or rather on $\tilde{\mathcal{T}_{S}}$, as both forms $T_{p^{\prime}}$ and $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ vanish along the line $\langle p^{\prime},p^{\prime\prime}\rangle$. By Lemma (3.4), this implies that $D_{L^{\prime},L^{\prime\prime}}$ divides $R_{i}$ as a function of $(p^{\prime},p^{\prime\prime})\in\tilde{\mathcal{T}_{S}}$. Eventually, the upshot is that $(D_{L^{\prime},L^{\prime\prime}})^{\mu}$ divides $R$, and since both $D_{L^{\prime},L^{\prime\prime}}$ and $R$ are well-defined as functions of $(p^{\prime},p^{\prime\prime})\in{\mathcal{T}_{S}}$, so is the quotient. Therefore, there exists a multi-homogeneous polynomial $R^{\prime}$ in the sets of variables $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$ such that after restriction to $\check{\mathbf{P}}^{3}\times\check{\mathbf{P}}^{3}\times\mathcal{T}_{S}$ one has $R=(D_{L^{\prime},L^{\prime\prime}})^{\mu}R^{\prime}$. Computing degrees, one finds that $R^{\prime}$ has multi-degree $(0,0,\mu(d-2)+\lambda,0)$ in $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$, which proves the theorem. ∎
@@
We recall that $\mathbf{k}$ is assumed to be an algebraically closed field of characteristic 0. We let $S\subseteq\mathbf{P}^{3}$ be a smooth surface of degree $d$, defined by a homogeneous polynomial $F(X,Y,Z,W)$. We consider generic $4$-tuples $\hat{p}^{\prime}=(X^{\prime},Y^{\prime},Z^{\prime},W^{\prime})$, $\hat{p}^{\prime\prime}=(X^{\prime\prime},Y^{\prime\prime},Z^{\prime\prime},W^{\prime\prime})$, $\hat{p}=(X,Y,Z,W)$, and call $p^{\prime},p^{\prime\prime},p$ the corresponding points in $\mathbf{P}^{3}$ (beware the unusual distribution of the prime decorations). The choice of $\hat{p}^{\prime},\hat{p}^{\prime\prime},\hat{p}$ defines a system of homogeneous coordinates $(\alpha:\beta:\gamma)$ on the plane $\langle p^{\prime},p^{\prime\prime},p\rangle$ generated by $^{\prime}, p′′p^{\prime\prime}$ and $p$. We consider
@@
The coefficient of $f(\alpha,\beta,\gamma)$ in $\alpha^{a}\beta^{b}\gamma^{c}$ is a polynomial of degree $a$ in the set of indeterminates $\hat{p}^{\prime}$, as can be seen on the Taylor–Newton expansion, the first terms of which are given in Display (3.1.1). Therefore, by (2.6.1), the discriminant $\mathrm{Disc}(f)$ is homogeneous of degree $d(d-1)^{2}$ in $\hat{p}^{\prime}$, and the dual surface $S^{\vee}$ has degree $d^{\vee}=d(d-1)^{2}$ in $\smash{\check{\mathbf{P}}^{3}}$.
@@
We consider the discriminant of the plane curve $S\cap\langle p^{\prime},p^{\prime\prime},p\rangle$, i.e., the discriminant of $f\in A[\alpha,\beta,\gamma]$ with $A=\mathbf{k}[X,Y,Z,W]$; by Formula (a), or rather Formula (2.1.2), it writes
@@
Salmon’s idea is that if the tangent plane $\mathbf{T}_{p^{\prime}}S$ is also tangent to $S$ at some point different from $p^{\prime}$, then it should appear with multiplicity >2>2 in $(p^{\prime},p^{\prime\prime})^{\perp}\cap S^{\vee}$; equivalently, $\mathrm{Disc}(f)$ should be divisible by $T^{3}$ in the above setup.
@@
The polynomial $A$ has degree 0 in $p$, therefore it is divisible by $T$ if and only if it is identically zero. By definition $A=\mathrm{D}^{\hat{p}^{\prime\prime 2}}f(\hat{p}^{\prime})$, so its vanishing is equivalent to the point $p^{\prime\prime}$ being on the polar quadric $\mathrm{D}_{{p^{\prime}}^{2}}S$ of $S$ at $p^{\prime}$. Since $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, this in turn is equivalent to the line $\langle p^{\prime},p^{\prime\prime}\rangle$ being one of the two inflectional tangents of $S$ at $p^{\prime}$ (i.e., the two lines $\Lambda$ having intersection multiplicity with $S$ at $p^{\prime}$ at least $3$; in symbols, $i(\langle p^{\prime},p^{\prime\prime}\rangle,S)_{p^{\prime}}\geqslant 3$).
@@
The polynomial $B^{2}-AC$ is tri-homogeneous of tri-degree (2(d2),2,2)(2(d-2),2,2) in the variables $(p^{\prime},p^{\prime\prime},p)$. By Theorem (3.2), arguing as in Paragraph (3.7) above, there exists a homogeneous polynomial $H$ of degree 4(d2)4(d-2) in the variable $p^{\prime}$, with constant coefficients, such that for fixed $p^{\prime}\in S$ and $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, $T$ divides $B^{2}-AC$ as polynomials in $p$ if and only if $H(p^{\prime})=0$.
@@
The discriminant in Display (3.4.2) is tri-homogeneous in the variables $(p^{\prime},p^{\prime\prime},p)$ with respect to the standard grading, of degree $d^{\vee}=d(d-1)^{2}$ with respect to all three variables, as we have observed above. This implies that also the homogeneous piece of (3.4.2) of lowest degree with respect to the Zariski grading is tri-homogeneous of tri-degree $(d^{\vee},d^{\vee},d^{\vee})$ for the standard grading. Moreover, $T$, AA, and $B^{2}-AC$ have respective tri-degrees (d1,0,1)(d-1,0,1), (d2,2,0)(d-2,2,0), and (2(d2),2,2)(2(d-2),2,2) in $(p^{\prime},p^{\prime\prime},p)$. Thus, computing degrees, one finds that $\Phi$ has tri-degree $(\lambda,\mu,\mu)$, with
@@
Since $\Phi$ is $\operatorname{redDisc}_{d}^{2}(f)$ with respect to the indeterminate $\alpha$, and the coefficient of $f(\alpha,\beta,\gamma)$ in $\alpha^{a}\beta^{b}\gamma^{c}$ is a polynomial of degree $a$ in the indeterminates $p^{\prime}$, the fact that $\Phi$ is homogeneous in $p^{\prime}$ and its degree follow directly from (2.14.1); to wit, in our situation the degree in (2.14.1) reads
@@
The homogeneity and degrees in $p$ and $p^{\prime\prime}$ however do not directly follow from Proposition (2.14), as for instance the coefficients in $A\beta^{2}+2B\beta\gamma+C\gamma^{2}$ do not all have the same weights in $p$ or $p^{\prime\prime}$. It is noticeable that $T^{2}A(B^{2}-AC)^{2}$ is homogeneous in $p$ and $p^{\prime\prime}$ because the weights of $T^{2}$ and $A$ balance each other; this is a very elementary reason why $T$ “should” appear with a square in Formula (3.4.2). The degree of $\operatorname{redDisc}_{d}^{2}(f)$ with the weights as in (2.14.2) is
@@
It has to be recognized that the proof given above of the latter statement rests on the not fully justified fact that $T$ divides $\Phi=\operatorname{redDisc}(f)$ if and only if the tangent plane $\mathbf{T}_{p^{\prime}}S$ is also tangent to $S$ at some additional point. The “if” part is established by [II, Theorem **], which is a result due to Dimca [9], but what is really needed is the “only if” part. Evidence in favour of that assertion is provided by the irreducibility of the reduced discriminant $\Phi$ (which, as already mentioned, follows from the fact that the reduced discriminant is a toric discriminant; see also [4]).
@@
In fact, Salmon’s procedure works in arbitrary dimension, using the generalization (2.2.6) of formula (a). The arguments are direct generalizations, so we are going to be sketchy.
Let $V=\mathrm{V}(F)$ be a smooth hypersurface of degree $d$ in $\mathbf{P}^{n}$, and consider $n$ points $p_{1},\ldots,p_{n-1},p$ (or rather tuples $\hat{p}_{1},\ldots,\hat{p}_{n-1},\hat{p}\in\mathbf{k}^{n+1}$). The points $p_{1},\ldots,p_{n-1}$ (if in linear general position, i.e., if $\hat{p}_{1},\ldots,\hat{p}_{n-1}$ are linearly independent) define a pencil of hyperplanes in $\mathbf{P}^{n}$, and for all $p$ (not in $\langle p_{1},\ldots,p_{n-1}\rangle$), the tuples $\hat{p}_{1},\ldots,\hat{p}_{n-1},\hat{p}$ define a system of homogeneous coordinates $(\alpha_{1}:\ldots:\alpha_{n-1}:\alpha)$ on the member $H_{p}$ of the pencil determined by $p$, i.e., $\langle p_{1},\ldots,p_{n-1},p\rangle$. In this system of coordinates, the hyperplane section $H_{p}\cap V$ is the zero locus of the polynomial
@@
where $T=\mathrm{D}_{\hat{p}}f(\hat{p}_{1})$ and $f_{2}(\alpha_{2},\ldots,\alpha_{n},\alpha)=\mathrm{D}^{(\alpha_{2}\hat{p}_{2}+\cdots+\alpha\hat{p})^{2}}F(\hat{p}_{1})$. By Formula (2.2.6), the generalization of (a), one has
@@
with $\bar{f}_{2}(\alpha_{2},\ldots,\alpha_{n})=f_{2}(\alpha_{2},\ldots,\alpha_{n},0)$ and $\Phi=\operatorname{redDisc}_{d}^{2}(f)$. Our task is to analyze the divisibilities of $\mathrm{Disc}(f_{2})$ and $\Phi$ by $T$ ($\bar{f}_{2}$ has degree 0 in $p$, hence $T$ divides $\mathrm{Disc}(\bar{f}_{2})$ if and only if $\mathrm{Disc}(\bar{f}_{2})=0$).
Theorem (3.2) (Salmon’s elimination trick) generalizes as follows. Given n1n-1 linear forms $L_{1},\ldots,L_{n-1}\in\mathbf{P}^{n}$, the line $\mathrm{V}(L_{1},\ldots,L_{n-1})$ intersects the (n2)(n-2)-dimensional linear space $\langle p_{1},\ldots,p_{n-1}\rangle$ if and only if
$D_{L_{1},\ldots,L_{n-1}}(p_{1},\ldots,p_{n-1})\overset{\text{def.}}{=}\det\bigl(L_{i}(p_{j})\bigr)_{1\leqslant i,j\leqslant n-1}$
@@
and consider a multihomogeneous polynomial $\Phi_{p_{1},\ldots,p_{n-1}}(p)$ of multidegree $(\lambda,\mu,\ldots,\mu)$ in the sets of variables $p_{1},\ldots,p_{n-1},p$, such that for all $(p_{1},\ldots,p_{n-1})\in\mathcal{T}_{V}$, the hypersurface $\mathrm{V}(\Phi_{p_{1},\ldots,p_{n-1}})$ consists of $\mu$ hyperplanes, counted with multiplicities, all containing the (n2)(n-2)-plane $\langle p_{1},\ldots,p_{n-1}\rangle$.
@@
in $p_{1}$, and $p_{2},\ldots,p_{n-1},p$, respectively (note that the former degree is divisible by d2d-2). As in Paragraph (3.9), the degree in $p_{1}$ of $\operatorname{redDisc}_{d}^{2}(f)$ is given by (2.14.1), while the degree with respect to the grading of (2.14.2) is equally distributed between the degrees in $p_{2},\ldots,p_{n-1},p$ respectively.
@@
Let $V$ be a smooth degree $d$ hypersurface in $\mathbf{P}^{n}$, $n>1$. The number of hyperplanes bitangent to $V$ passing through n2n-2 fixed general points in $\mathbf{P}^{n}$ is
@@
In this section, we elaborate on the computation by Salmon of the number of bitangent lines to a surface in $\mathbf{P}^{3}$ passing through a general point. We give his proof in Section 4.1, and a generalization to hypersurfaces in $\mathbf{P}^{n}$ in the next Section 4.2. These proofs only involve the reduced discriminant of polynomials in two indeterminates, which is easily expressed as a plain discriminant, see Example (2.10). We give however another generalization in Section 4.3, for which it is necessary to consider reduced discriminants of polynomials with an arbitrary number of indeterminates. The two statements of Sections 4.1 and 4.2 are instances of a vast generalization by Fehér and Juhász [10], which is described in [III, Theorem **].
@@
There exists a polynomial $R$, bihomogeneous in $p^{\prime},p^{\prime\prime}$ with respective degrees (d2)(d3)(d-2)(d-3) and (d+2)(d3)(d+2)(d-3), such that the locus of points $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$ for which the line $\langle p^{\prime},p^{\prime\prime}\rangle$ is bitangent to $S$ is cut out on $\mathcal{T}_{S}$ by $V(R)$.
@@
(this is an abuse of notation: actually one should consider two liftings $\hat{p}^{\prime},\hat{p}^{\prime\prime}\in\mathbf{k}^{4}$ of $p^{\prime}$ and $p^{\prime\prime}$ respectively, and $F(\alpha\hat{p}^{\prime}+\beta\hat{p}^{\prime\prime})$); the last equality comes from the fact that $\langle p^{\prime},p^{\prime\prime}\rangle$ is tangent to $S$ at $p^{\prime}$. Let $\vphantom{f}{}^{\flat}\kern-1.00006ptf(\alpha,\beta)$ be the homogeneous polynomial of degree d2d-2 between parentheses at the right-hand-side of (4.2.1). The line $\langle p^{\prime},p^{\prime\prime}\rangle$ is a bitangent to $S$ if and only if $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ has a multiple root, so the polynomial $R$ we are looking for is merely the discriminant of $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ which is none other, up to sign, than the reduced discriminant of $f$, as noticed in (2.10).